Archive for April 2016

Next Meeting: Sunday 1st May Looking good for full inspection.

Forecast look ok for full inspection , if done quickly.  Main thing is to check if queen right, ie, new eggs, or seeing the queen, checking  room,checking  stores, and any signs of disease.  It was noticed at last weeks inspection that several colonies had a few chalkbrood on the varroa board ( vb) .   This could have been caused by the low temperatures the week before when we inspected the colonies but something to keep an eye on.

Looking forward to a busy Sunday afternoon.   If you can arrive about 14:00 to get kitted up asap for a 14:15 start that would be great and don’t forget to wear wellies or something similar to enable dunking in sterilizing solution before entering the apiary.

See you there.

Richard, Liz and Mel.

Apiary Report: 24th April

Unfortunately the weather forecast was correct and it was only 8 – 10 C, and overcast, but the sun did come out for a time, but still too cold to go through colonies, still we had other jobs to do, checking varroa boards, checking rapid feeders put on last week , burning up some of the winter debris and branding super boxes with the club stamp.

We had nine members helping out and as usual finished with a lovely cup of tea and cakes, thanks Pippa and Liz, along with lots of discussions on the state of our bees and the effect the cold weather is having on them.

Varroa counts were high on the colonies treated with icing sugar, showing that it had some effect.  We will try again next time we open up colonies.

Varroa count over 7 days

Hive#1      1     some chalk brood and mouse droppings on vb

Hive#2     40 ( icing sugar last week)

Hive#3      0    topped up feeder

Hive#4     23

Hive#5     25    mouse droppings on vb

Hive#6      4     Chalk brood and mouse droppings on vb

Hive#7     30    (Icing Sugar last week)   chalk brood on vb

Hive#8     8     Chalk brood on vb

Hive#9   Warre –  no OMF.

Hive#10   0    syrup still available, small number of bees.

Hive#15F    1    Nuc still going strong

Hive#15H   7    Poly nuc also still active

Apiary Report : 17th April

The forecast for Sunday was 10 C but luck was on our side and we had 12 -14 C and lovely sunshine,  so it was decided to carry on and go through the colonies for the first time, with special attention to number 7 as it had a very high ( 50 in a week) varroa count.   As a first measure we will try dusting with icing sugar to encourage bees to groom themselves and clear mites at the same time.

Temperature :  12 -14 C,    Sunny,  little wind.

Hive#1   Inspected by Brian Herbert

Brood box :   14 x 12 ,    over 3 frames of brood at all stages, stores ok, room ok, varroa count 4.  temper 10.

Queen excluder and super added,  some pepper-potting with drone cells needs an eye kept on. may mean the queen is losing her fertility.

Hive#2    Inspected  by Brian Herbert

Brood box:  14 x 12  and half,   1 frame in Deep, 2 frames of brood in half.  stores, room ok.  Queen excluder added , under super put back on top.  varroa count 30,  so dusted with icing sugar.   Temper 10

Hive#3    Inspected by Mel.

Brood box:  Deep National.  Small colony, only half a frame of brood, no stores, needed feeding 50:50 sugar solution added.   varroa count  0,   Temper 10

Hive#4   Inspected by Mel.

Brood box:  Deep National,  4 frames of brood, room ok, no stores, needs feeding. 50:50 sugar solution added.    Qx added under super moved from under brood box. Varroa count 12,   Temper 10.

Hive#5   Inspected by Mel.

Brood box:   Deep National,   3 frames of brood.  room ok,  no stores, 50:50 added.  Qx added under super, moved up from under brood box.  Drone brood and 1 drone seen, 3 mouldy super frames swapped.

Hive#6   Inspected by  Liz.

Brood box:  Deep National.  Brood all stages (BAS)  Q seen -still blue. stores and room ok. varroa count 3.  Temper 8 mouse dropping again on vb.

Hive#7   Inspected by Liz.

Brood box:    All Halves (supers)  Peter Streets Method.    6 -7 frames of B A S.  stores ok, room ok.  Big colony.  Varroa count 30 but last week was 50, dusted frame tops and bees with icing sugar.   also a few chalkbrood.

Hive#8   Inspected by Liz.

Brood box:  National Deep.  Q seen,still blue.  brood ok, stores and room ok.  Under super moved up above Qx and another super added.  varroa count 5.   Temper 9

Hive#9   Warre Hive   Quick check only, light stores so 50:50  feed put on.

Hive#10  Inspected by Richard/Brian

Top Bar Hive,   one extra bar put in main chamber,  small colony but stores and room ok. 50:50 syrup put on to help.

Hive 15 F   Inspected by Liz.

Brood box :   National Deep.  Q seen, still blue.  B A S.  stores and room ok.    varroa count  3,  Temper 8

Hive 15 H   Inspected by Liz.

Poly Nuc.  Making up nicely.   1 drawn frame moved next to brood frames.

 

Richard, Liz and Mel.  Apiary Custodians

Next Meeting : 24th April – Sunday meeting still on but no inspection of hives.

Low temperatures again forecast but we have things to do left from last week.     Sorry all, but it is just too cold to go through hives.

Learn how to assemble flat pack supers,  frame cleaning,  clearing and tidying up area with bonfire if wind direction away from apiary.  checking varroa boards especially no 7 to see if icing sugar has had any effect.

 

See you all there.

Richard, Liz and Mel.     Apiary custodians

Next Apiary Meeting: 17th April still on but no inspections

Again we have possible problems with the weather being too cold for a full inspection of colonies but there are things we can be doing.

One colony  no7  had a high varroa count  so we will be treating it with icing sugar to see if has any effect.

We can make up a couple of super boxes.

Clear up surrounding apiary area of dead branches and have a bonfire if wind direction ok.

Sorry all you keen beekeepers but we have to be patient and wait till the conditions are right.

Richard, Liz and Mel.       Apiary custodians

Apiary Meeting: 10th April Cancelled due to cold weather

The forecast for Sunday is again very marginal for checking bees ,  so its best to wait till warmer times before a full inspection.  Sorry to disappoint all you new keen  beekeepers and the older members who would like a refresh on handling bees.   Lets hope next week will be better.

Richard, Liz and Mel.   Apiary Custodians.

Apiary: April 3rd 2016

General check, mouse guard removal and replacement with restricted entrances, stores check and varroa check.    Weather  Cloudy,  12-14 C  wind  f3 SE

All colonies had bees flying with pollen coming in.

Hive#1 Mouse droppings on varroa board ( vb)  old hard fondant removed and replaced , 17 v in 21 days.

Hive#2  Old fondant removed,  17 v in 21 days

Hive#3  Old fondant removed and new   put on.     5 v in 21 days

Hive#4 Old fondant removed     No varroa

Hive#5 Old fondant removed,    5 v in 21 days

Hive#6 Old fondant removed and new put on. mouse droppings on vb.   No varroa

Hive#7 Old fondant removed,    25 v in 21 days

Hive#8 Old fondant removed ,  3  v in 21 days

Hive#9 Warre hive.  Old fondant removed . unable to check for v

Hive#10 Top bar hive,  Old fondant removed,   mouse droppings on top bars.  No varroa

Hive 15F  Nuc.  Old fondant removed.   5 v in 21 days

Hive 15H  Poly Nuc,  Old fondant removed.   22 v in 21 days.

Apiary: First Meeting April 3rd

The weather was a lot colder than forecast, cloudy with a brisk southeasterly breeze making it too cold to go through the colonies but there were some things we could do.

The mouse guards were removed and replaced with restricted entrance blocks.

Any remaining fondant and clingfilm left was removed as it had gone solid and difficult for the bees to use.  Some colonies  that appeared low on stores were resupplied with new fondant patties.

Varroa boards were checked,cleaned and replaced.

The twelve new members that attended were shown the various hives in use and the advantages and disadvantages of each method of beekeeping.

All in all a useful run through ready for the better weather to come and a full inspection of the colonies.

If the weather is doubtful on a Sunday please check the website to see if the meeting is to go ahead.

Richard , Liz and Mel.